OpenStudy (anonymous):

Solve : -4(x+10)-6=-3(x-2)

OpenStudy (anonymous):

Um, so...

OpenStudy (anonymous):

x=54

OpenStudy (anonymous):

-4x-40-6 =-3x+6 or -4x-46=-3x+6 or -46-6 =-3x+4x or -54 = x